OpenStudy (anonymous):

solve by factoring: 5x^2 + 9x - 72 = 0

OpenStudy (anonymous):

@jim_thompson5910 can you help? I just cant figure out the factors that would sum up to 9

jimthompson5910 (jim_thompson5910):

this might be a bit ugly

jimthompson5910 (jim_thompson5910):

5*(-72) = -360

jimthompson5910 (jim_thompson5910):

you need to find two numbers that a) multiply to -360 AND b) add to 9

OpenStudy (anonymous):

15 & 24!

jimthompson5910 (jim_thompson5910):

so you would list out the ways to multiply to -360 (ie factor it) -1*360 -2*180 -3*120 ... ... etc etc then add up each factor pair -1+360 = 359 ...is that equal to 9? no, so move on -2+180 = 178...is that equal to 9? no, so move on -3+120 = 117...is that equal to 9? no, so move on etc etc

jimthompson5910 (jim_thompson5910):

15+24 = 39 not 9

jimthompson5910 (jim_thompson5910):

however 24 + (-15) = 9

jimthompson5910 (jim_thompson5910):

5x^2 + 9x - 72 5x^2 + 24x - 15x - 72 now factor by grouping

OpenStudy (anonymous):

5x^2 + 9x - 72 = 0 I suppose we could use the quadratic formula? \[\frac{ -9\pm \sqrt{9^{2}-4(5)(-72)} }{ 2(5) }\]

OpenStudy (anonymous):

Wouldn't it be... (5x + 24)(x - 15)

OpenStudy (anonymous):

I mean we can find the zeroes and then just group them?

OpenStudy (anonymous):

5x + 24 = 0 5x = -24 x = -24/5 x - 15 = 0 x = 15

jimthompson5910 (jim_thompson5910):

yes you could use the quadratic formula to find the zeros/roots then you use that to find the factorization that's probably the most efficient method

OpenStudy (anonymous):

Exactly.

jimthompson5910 (jim_thompson5910):

but here's how you factor it by grouping 5x^2+9x-72 5x^2+24x-15x-72 (5x^2+24x)+(-15x-72) x(5x+24)+(-15x-72) x(5x+24)-3(5x+24) (x-3)(5x+24) --------------------------------------------------------------------------- So 5x^2+9x-72 completely factors to (x-3)(5x+24)

OpenStudy (anonymous):

actually its x = 3 and x = -24/5

OpenStudy (anonymous):

^ @angelina22309

OpenStudy (anonymous):

ok I get it, How do you know when you need to factor by graphing? Why woulnt have mine worked?

OpenStudy (anonymous):

Could you also help with 3x^2 + 19x - 50 - 0

jimthompson5910 (jim_thompson5910):

your method only works when the leading coefficient is 1

jimthompson5910 (jim_thompson5910):

example x^2 + 5x + 6 factors to (x+3)(x+2) since 3+2 = 5 and 3*2 = 6

OpenStudy (anonymous):

oh ok. thanks!! I need to keep that in mind

OpenStudy (anonymous):

@jim_thompson5910 Please let me know if I did this correctly. 3x^2 + 19x - 50 = 0 3x^2 + 25x - 6x - 50 (3x^2 + 25x) + (-6x - 50) x(3x + 25) + (-6x - 50) x(3x + 25) -2(4x + 25) (x - 2)(4x + 25)

OpenStudy (anonymous):

@some_someone could you check this please ^

jimthompson5910 (jim_thompson5910):

3x^2+19x-50 3x^2+25x-6x-50 (3x^2+25x)+(-6x-50) x(3x+25)+(-6x-50) x(3x+25)-2(3x+25) (x-2)(3x+25) --------------------------------------------------------------------------- 3x^2+19x-50 completely factors to (x-2)(3x+25)

jimthompson5910 (jim_thompson5910):

so the only mistake you made was writing 4x instead of 3x on this line x(3x + 25) -2(4x + 25)

OpenStudy (anonymous):

thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw